OpenStudy (anonymous):

How do I write y=x^2+4 and y=x^2+14x+50 in vertex form?

OpenStudy (anonymous):

The vertex form of the given quadratic f(x) = x^2 - 14x + 50 is f(x) = (x - 7)^2 + 1.

OpenStudy (anonymous):

does that help

OpenStudy (precal):

you will need to complete the square